# Enhancing Startup Success Predictions in Venture Capital: A GraphRAG Augmented Multivariate Time Series Method

### v1

@@ Introduction
Consider an undirected graph $G(V,E,w)$, where $V$ represents the set of vertices, $E$ the set of edges, and $w_{ij}=w_{ji}$ denotes the weight associated with each edge. In the case of an unweighted graph, we assume unit weight for each edge ($w_{ij}=1$). Additionally, the neighbors of a vertex $i$ are denoted as $J_{i}=\{j\ |\ (i,j)\in E\}$, the weighted degree of each vertex as $K_{i}=\sum_{j\in J_{i}}w_{ij}$, the total number of vertices as $N=|V|$, the total number of edges as $M=|E|$, and the sum of edge weights in the undirected graph as $m=\sum_{i,j\in V}w_{ij}/2$.
Disjoint community detection involves identifying a community membership mapping, $C:V\rightarrow\Gamma$, where each vertex $i\in V$ is assigned a community-id $c$ from the set of community-ids $\Gamma$. We denote the vertices of a community $c\in\Gamma$ as $V_{c}$, and the community that a vertex $i$ belongs to as $C_{i}$. Further, we denote the neighbors of vertex $i$ belonging to a community $c$ as $J_{i\rightarrow c}=\{j\ |\ j\in J_{i}\ and\ C_{j}=c\}$, the sum of those edge weights as $K_{i\rightarrow c}=\sum_{j\in J_{i\rightarrow c}}w_{ij}$, the sum of weights of edges within a community $c$ as $\sigma_{c}=\sum_{(i,j)\in E\ and\ C_{i}=C_{j}=c}w_{ij}$, and the total edge weight of a community $c$ as $\Sigma_{c}=\sum_{(i,j)\in E\ and\ C_{i}=c}w_{ij}$ (Leskovec 2021).
Modularity serves as a fitness metric for evaluating the quality of communities identified by heuristic-based community detection algorithms. It is calculated as the difference between the fraction of edges within communities and the expected fraction if edges were randomly distributed, yielding a range of $[-0.5,1]$, where higher values signify superior results (Brandes et al. 2007).The optimization of this metric theoretically leads to the optimal grouping (Newman 2004). The modularity $Q$ of identified communities is determined using modularity equation, where $\delta$ represents the Kronecker delta function ($\delta(x,y)=1$ if $x=y$, 00 otherwise). The delta modularity of moving a vertex $i$ from community $d$ to community $c$, denoted as $\Delta Q_{i:d\rightarrow c}$, can be calculated as follows:

### v2

@@ Introduction
Let $G(V,E,w)$ be an undirected graph where $V$ represents the vertex set, $E$ represents the edge set, and $w_{ij}=w_{ji}$ represents the weight of the edge between vertices $i$ and $j$. For unweighted graphs, each edge is assumed to have a unit weight, i.e., $w_{ij}=1$. The set of neighbors for a vertex $i$ is defined as $J_{i}=\{j\ |\ (i,j)\in E\}$. The weighted degree of a vertex $i$ is given by $K_{i}=\sum_{j\in J_{i}}w_{ij}$. The total number of vertices is denoted by $N=|V|$, the total number of edges by $M=|E|$, and the sum of all edge weights in the graph is $m=\sum_{i,j\in V}w_{ij}/2$.
Disjoint community detection focuses on determining a mapping of community membership, represented as $C:V\rightarrow\Gamma$, where each vertex $i\in V$ is associated with a community identifier $c$ from the set of community identifiers $\Gamma$. The vertices belonging to a community $c\in\Gamma$ are denoted by $V_{c}$, and the community to which a vertex $i$ belongs is represented by $C_{i}$. Additionally, the neighbors of vertex $i$ that are part of community $c$ are denoted as $J_{i\rightarrow c}=\{j\ |\ j\in J_{i}\ and\ C_{j}=c\}$, with the total weight of these edges given by $K_{i\rightarrow c}=\sum_{j\in J_{i\rightarrow c}}w_{ij}$. The total weight of edges within community $c$ is represented by $\sigma_{c}=\sum_{(i,j)\in E\ and\ C_{i}=C_{j}=c}w_{ij}$, and the overall edge weight associated with community $c$ is denoted by $\Sigma_{c}=\sum_{(i,j)\in E\ and\ C_{i}=c}w_{ij}$ (Leskovec 2021).
Modularity is a measure used to assess the quality of communities identified by heuristic-based community detection algorithms. It is determined by calculating the difference between the actual fraction of edges within communities and the expected fraction of such edges if they were randomly distributed. The resulting value typically falls within the range of $[-0.5,1]$, where higher values signify superior results (Brandes et al. 2007). The optimization of this metric theoretically leads to the optimal grouping (Newman 2004). The modularity $Q$ of identified communities is determined using modularity equation, where delta represents the Kronecker delta function ($\delta(x,y)=1$ if $x=y$, 00 otherwise). The delta modularity of moving a vertex $i$ from community $d$ to community $c$, denoted as $\Delta Q_{i:d\rightarrow c}$, can be calculated as follows: